# Advances in Appfl: A Comprehensive and Extensible Federated Learning Framework

### v1

@@ III-C Server Scheduling and Aggregation
In order to tackle the computation heterogeneity in FL where clients have varying computing capabilities, many asynchronous aggregation algorithms have been proposed to reduce client idle times and enhance resource utilization. To support aggregation with different synchronicity settings, Appfl introduces a server-side scheduler that acts as an interface between the communicator and the aggregator. Upon receiving a local model from a client, the communicator forwards it to the scheduler, which determines the appropriate time to pass the local model(s) to the aggregator for updating the global model. For synchronous aggregation strategies, such as FedAvg [3], a synchronous scheduler buffers each client’s local model until all models are received, at which point it forwards them to the aggregator to update the global model. Conversely, for asynchronous strategies like FedAsync [36], a vanilla asynchronous scheduler immediately sends the client model to the aggregator and returns the updated global model back to the communicator. Additionally, the scheduler module is designed to be extensible for the incorporation of more advanced scheduling algorithms. Specifically, Appfl supports the state-of-the-art Compass asynchronous scheduler [23], aiming to alleviate the drift of the global model toward faster clients. Such drift is prevalent in other asynchronous FL algorithms, where faster clients update the global model more frequently and the models from slower clients become stale. The Compass scheduler synchronizes the arrival of a group of client local models by assigning different amounts of local training tasks to different clients to enable a grouped global aggregation and avoid stale local models, mitigating the client drift issue. Figure 3 illustrates the scheduling processes under the three different schedulers.
As for the aggregator module, Appfl supports a broad range of aggregation strategies, going beyond the widely used FedAvg. These include FedAvgM, FedAdam, and FedYogi, which address data heterogeneity, PLFL [88] for personalized FL, as well as ICEADMM [89] and IIADMM [22], which focus on efficient privacy preservation. Additionally, for asynchronous aggregation, Appfl includes strategies such as FedAsync, FedBuff, AREA [39], and FedCompass. This diverse suite of options ensures that Appfl can accommodate a variety of needs and scenarios in FL, illustrating its adaptability and comprehensive approach to FL challenges.
@@ III-D Privacy Preservation and Authentication
To tackle security and privacy concerns in FL, Appfl offers solutions that span both algorithmic and system-level measures. Algorithmically, Appfl incorporates the differential privacy (DP) algorithms [30] into the FL process that perturbs the client model parameters with noises before sending to the server, protecting against the reconstruction of confidential training data. A study utilizing Appfl showcases that the usage of DP in FL can effectively mitigate the risk of data reconstruction [8].
@@
As for gRPC, Appfl utilizes token-based authenticators to verify users. Clients have to attach an access token to each remote procedure call (RPC) request over an SSL-encrypted channel, allowing the server to confirm the user’s identity before processing the request. The token-based authenticator consists of two primary functions: one invoked by the client to generate the token prior to sending the RPC request and another invoked by the server to verify the validity of the token upon receipt. This straightforward interface allows developers to effortlessly integrate their own authentication methods tailored to specific use cases and applications. Currently, Appfl supports a Globus authenticator, with its login flow depicted in Figure 4. Users can employ Appfl’s command line interface (CLI), appfl-auth, to perform a one-time login. Depending on the selected role during login, either as an FL server or client, the appropriate Globus access token (Group Service or Identity Service) is requested. The access tokens, along with the corresponding refresh tokens, are securely stored in the client’s local token storage. Whenever an FL client makes an RPC request, it attaches its Globus Identity Service token. The FL server uses this token to retrieve the client’s Globus ID and, leveraging its Globus Group Service token, verifies whether the client belongs to the specified Globus group. This robust authentication process ensures a secure and controlled federation for FL experiments.
@@ IV Performance Evaluation
In this section, we employ Appfl to comprehensively benchmark a broad spectrum of aspects within FL to highlight the capabilities of its new software design. Specifically, we utilize Appfl to evaluate the communication efficiency of different communication protocols, data transfer methods, and compression algorithms. We also explore the impacts of privacy preservation algorithms on the performance of FL-trained models, as well as the training efficiency and resource utilization of various FL strategies under different synchronicity settings. These comprehensive assessments highlight the capabilities of the newly designed Appfl framework in benchmarking a broad spectrum of aspects within FL.
@@ IV-A Communication Efficiency
We evaluate the communication efficiency for different communication and data transfer protocols across different numbers of clients and various model sizes, ranging from a few-byte 1$\times$1 fully connected (FC) layer to a Vision Transformer (ViT) with hundreds of megabytes. Table II details the sizes of all models used in our experiments.
@@
Model Params Size
@@
CNN [90] 1.20M 4.58 MB
ResNet18 [91] 11.17M 42.66 MB
ResNet50 [91] 23.52M 89.93 MB
ResNet101 [91] 42.51M 162.58 MB
Vision Transformer [92] 88.22M 336.55 MB
In the experiments, each FL client runs on a single core of the CPU-only compute nodes of the Delta supercomputer at the National Center for Supercomputing Applications. Each CPU node contains two 64-core AMD EPYC 7763 “Milan” CPUs with PCIe Gen4 interfaces and 256 GB of RAM. Delta is connected to the NPCF core router and exit infrastructure via two 100 gigabits per second (Gbps) connections. The FL server is hosted on an AWS EC2 x2iedn.2xlarge instance, equipped with 8 virtual CPUs, 256 GB of RAM, and up to 25 Gbps connections. We exponentially increase the number of clients from 2 to 128 across all models, except for the ViT model, which scales only from 2 to 64 because of memory constraints on the client and server hardware. We evaluate gRPC and Globus Compute communication protocols as well as two data transfer methods, AWS S3 buckets and ProxyStore endpoints. Because of the 10 MB data transfer limit with Globus Compute, it is integrated with other data transfer protocols rather than being tested in isolation, resulting in five distinct communication pattern combinations.
Figure 5 shows the epoch-wise two-way communication time in seconds for various models using different communication and data transfer protocols. From the plots, we note the following key points: (1) Separating the transmission of data (i.e. model parameters) from task control signals helps communication protocols exceed their maximum data size limitations. (2) Globus Compute consistently incurs longer overheads than does gRPC in transmitting control signals, which is a significant factor when the FL model size is small. (3) While data transfer via ProxyStore endpoints generally results in longer communication times, it offers a free and straightforward solution for protocols such as Globus Compute that have message size restrictions. (4) Data transfer through S3 features relatively low latency and also provides a secure, reliable means to store model checkpoints during training, although it incurs some costs.
@@ IV-B Compression Efficiency
We assess the efficiency of various data compression algorithms integrated within the Appfl framework. Specifically, we utilize the lossless compressor blosc [84] for tensors with less than 1,024 parameters and lossy compressors SZ2 [85], SZ3 [86], and ZFP [87], each with a relative error bound of 0.01, for larger tensors. The experiments adhere to the same hardware configurations described in Subsection IV-A, with clients running on Delta and the server on an AWS EC2 instance. We conduct experiments on ResNet-50 and ResNet-101 models, scaling client numbers from 2 to 128. Figure 6(a) illustrates the reduction in model sizes by 3 to 5 times using different lossy compressors. Notably, previous studies have shown that such levels of lossy compression can preserve model accuracy within a 0.5% margin of uncompressed results [26]. Figure 6(b) presents the two-way communication times via gRPC for the two models using various compressors. Solid lines represent times with compression and decompression overheads, whereas dotted lines depict times without. The comparison reveals significant overhead, particularly with SZ2 and SZ3. Despite this, the use of compressors notably reduces communication costs and overall two-way communication times, even under high-bandwidth conditions for both clients and the server.
@@ IV-C Privacy Preservation
Model Cox model [94] Logistic Reg. 3D U-Net [95] EfficientNet [96]
Metric C-index Accuracy DICE [97] Balanced Acc.
@@ IV-D Addressing Heterogeneous Clients
Figure 8: (a) Average validation accuracy and the corresponding standard deviation on the partitioned CIFAR-10 dataset for different FL algorithms during the training process. (b) Client resource utilization for algorithms using different schedulers and the average training time per batch for different clients. (c) Visualization of computing resource utilization for three clients under different schedulers, where colored bar represents the computing period and the blank places means idle times.
@@
In this subsection we evaluate the performance and efficiency of different FL algorithms under various synchronicity settings. Specifically, we benchmark five FL algorithms: (1) FedAvg, a widely used synchronous algorithm that updates the global model by averaging all client local models; (2) FedAvgM, another synchronous algorithm, which incorporates momentum on top of FedAvg; (3) FedAsync, which asynchronously updates the global model upon receipt of any local model; (4) FedBuff, which is similar to FedAsync but buffers multiple local models before updating the global model; and (5) FedCompass, which introduces a COMputing Power AwarenesS Scheduler (Compass) that dynamically adjusts the number of client local training steps based on real-time estimates of client computing power to synchronize the training completion for groups of clients. As for the datasets, we partition the CIFAR-10 dataset [98], one of the most commonly used datasets in evaluating FL algorithms [99], into ten client splits in a non-IID manner, with each client holding data from five to seven classes out of ten classes. All clients use Nvidia A100 GPUs for training, and we simulate a group of heterogeneous clients by assigning different average batch processing times from an exponential distribution.
@@
Figure 8(a) presents the average validation accuracy and the corresponding standard deviation across five independent runs for each FL algorithm during training. Key observations from the figure include the following. (1) Asynchronous FL algorithms like FedAsync and FedBuff, which use the vanilla asynchronous scheduler, converge to significantly lower global model accuracy compared with synchronous methods, primarily due to the drifting toward faster clients, as the global model gets more updates from faster clients and slower clients’ models become stale. (2) Synchronous algorithms exhibit slower convergence as the server has to wait for the slow clients for aggregation. (3) FedCompass effectively addresses substantial client drift issues and attains high global model accuracy by ensuring nearly simultaneous model arrivals for grouped aggregation. It also achieves quicker convergence than synchronous methods without extensive waiting.
@@
Figure 8(b) shows the average training time per batch for the ten clients involved in the FL training, as well as the resource utilization, calculated as the ratio of client compute time to total training time, for algorithms using the synchronous, vanilla asynchronous, and Compass asynchronous scheduler. The synchronous scheduler shows the lowest client resource utilization, correlating with training time per batch: the quicker the client, the lower the rutilization. In contrast, the vanilla asynchronous scheduler, which immediately sends any received local model for aggregation and returns the updated global model, allows client resource utilization to approach 100%. Despite full utilization, however, this method results in poorly performing models due to client drift. The Compass scheduler, by estimating client speeds and adjusting training steps accordingly, maintains approximately 90% resource utilization and reduces client drift through timely grouped aggregations. Figure 8(c) visualizes the resource utilization for three clients under different scheduling scenarios, highlighting the significant resource underutilization of the synchronous scheduler compared with the asynchronous alternatives when client computing resources vary widely.
@@ V-A Vertical Federated Learning
Vertical federated learning (VFL) is a specialized paradigm of FL where different clients hold distinct features from the same dataset [79]. Unlike traditional FL (i.e., horizontal FL) dealing with the same feature space across diverse data samples, VFL enables collaboration among clients that have partially overlapping or non-overlapping features but share the same sample IDs, as illustrated in Figure 9(a). Figure 9(b) depicts a typical VFL process. In VFL, rather than training the same model architecture and sharing model parameters, each client possesses its embedding model to process its local data sample features and then sends their embeddings to the server. The server, holding the labels of the client data samples, concatenates the received embeddings to train a central model. It then sends the gradients of the feature embeddings back to the corresponding clients, enabling them to update their local embedding models accordingly.
Appfl seamlessly supports VFL by providing the VFL trainer and aggregator in the corresponding modules. In this case study, we use the diabetes datasets [100] from the scikit-learn library, which contains ten features of 442 data samples. The labels, ranging from 25 to 346, are the responses of interest that quantitatively measure the disease progression. We split the dataset into 80% for training and 20% for validation and use three VFL clients, where clients 1 and 2 possess three patient features and client 3 possesses four. Each of the three clients as well as the server employs a two-layer perceptron with ReLU nonlinear activation [101] as their embedding models. Figure 10(a) presents the input, hidden, and output dimensions of these models. During the training, the server model is updated based on the mean squared error (MSE) loss between the labels and predictions, using the Adam optimizer [102] with a learning rate of 0.01. Figure 10(b) shows the training and validation MSE throughout the training.
@@ V-B Hierarchical Federated Learning
Figure 11: (a) Topology of the multi-layer HierFL experiments with nine clients, five intermediate servers, and one root server. (b) HierFL validation accuracy for the server and client models, where the accuracy of client models is evaluated after each local training step.
Hierarchical federated learning (HierFL) is also a special type of FL that introduces an additional role, the intermediate server (edge server). This server first aggregates local model parameters from connected clients or child intermediate servers and then forwards the aggregated model to the parent server for further aggregation [80]. HierFL is particularly beneficial when FL clients are geographically clustered, since placing an intermediate server for these clusters can significantly improve overall communication efficiency. To support HierFL in Appfl, in addition to the general server agent for the root server and the client agent for the clients, we define an intermediate server agent, similar to the server agent, which handles FL-related requests from connected clients or child intermediate servers by interacting with its parent server.
In this case study we conduct four-tier HierFL experiments involving nine clients, five intermediate servers, and one root server. The MNIST training dataset [103] is partitioned into nine heterogeneous client splits, with each client containing training data for only three to five classes. Figure 11(a) illustrates the topology of the HierFL experiments and the client data distribution. Training is conducted over 20 global epochs, with each client performing 100 local steps per epoch using a batch size of 64 and the Adam optimizer with a learning rate of 0.001. The experiments are repeated five times with different random seeds. Figure 11(b) presents the average validation accuracy and standard deviation for both the server model and each client’s local model on the MNIST validation set. We note that the client models are evaluated after local training. Since each client has data for only three to five classes, their local models perform significantly worse than the global model, highlighting the advantages of federated learning in leveraging data from distributed clients to train a more robust ML model.

### v2

@@ III-C Server Scheduling and Aggregation
In order to tackle the computation heterogeneity in FL where clients have varying computing capabilities, many asynchronous aggregation algorithms have been proposed to reduce client idle times and enhance resource utilization. To support aggregation with different synchronicity settings, Appfl introduces a server-side scheduler that acts as an interface between the communicator and the aggregator. Upon receiving a local model from a client, the communicator forwards it to the scheduler, which determines the appropriate time to pass the local model(s) to the aggregator for updating the global model. For synchronous aggregation strategies, such as FedAvg [2], a synchronous scheduler buffers each client’s local model until all models are received, at which point it forwards them to the aggregator to update the global model. Conversely, for asynchronous strategies like FedAsync [10], a vanilla asynchronous scheduler immediately sends the client model to the aggregator and returns the updated global model back to the communicator. Additionally, the scheduler module is designed to be extensible for the incorporation of more advanced scheduling algorithms. Specifically, Appfl supports the state-of-the-art Compass asynchronous scheduler [16], aiming to alleviate the drift of the global model toward faster clients. Such drift is prevalent in other asynchronous FL algorithms, where faster clients update the global model more frequently and the models from slower clients become stale. The Compass scheduler synchronizes the arrival of a group of client local models by assigning different amounts of local training tasks to different clients to enable a grouped global aggregation and avoid stale local models, alleviating the client drift issue. The seamless integration of FedCompass further exemplifies our framework’s extensibility. Figure 3 illustrates the scheduling processes under the three different schedulers.
As for the aggregator module, Appfl supports a broad range of aggregation strategies, going beyond the widely used FedAvg. These include FedAvgM, FedAdam, FedAdagrad, and FedYogi, which address data heterogeneity, PLFL [53] for personalized FL, as well as IIADMM [15], which focuses on efficient privacy preservation. Additionally, for asynchronous aggregation, Appfl includes strategies such as FedAsync, FedBuff, and FedCompass. This diverse suite of options ensures that Appfl can accommodate a variety of needs and scenarios in FL, illustrating its adaptability and comprehensive approach to FL challenges.
@@ III-D Privacy Preservation and Authentication
To tackle security concerns in FL, Appfl offers solutions that span both algorithmic and system-level measures. Algorithmically, Appfl incorporates differential privacy (DP) algorithms [18] into FL that perturbs the client model parameters with noises before sending to the server, protecting against the reconstruction of confidential training data. A study utilizing Appfl showcases that the usage of DP in FL can effectively mitigate the risk of data reconstruction [5].
@@
As for gRPC, Appfl utilizes token-based authenticators to verify users. Clients have to attach an access token to each remote procedure call (RPC) request over an SSL-encrypted channel, allowing the server to confirm the user’s identity before processing the request. The token-based authenticator consists of two primary functions: one invoked by the client to generate the token prior to sending the RPC request and another invoked by the server to verify the validity of the token upon receipt. This straightforward interface allows developers to effortlessly integrate their own authentication methods tailored to specific use cases and applications. Currently, Appfl supports a Globus authenticator, with its login flow depicted in Figure 4. Users can employ Appfl’s command line interface (CLI), appfl-auth, to perform a one-time login. Depending on the selected role during login, either as an FL server or client, the appropriate Globus access token (Group Service or Identity Service) is requested. The access tokens, along with the corresponding refresh tokens, are securely stored in the client’s local token storage. Whenever an FL client makes an RPC request, it attaches its Globus Identity Service token. The FL server uses this token to retrieve the client’s Globus ID and, leveraging its Globus Group Service token, verifies whether the client belongs to the specified Globus group. This robust authentication process ensures a secure and controlled federation for FL experiments, and significantly reduces the risk of malicious attack.
@@ IV Performance Evaluation
In this section, we employ Appfl to benchmark a broad spectrum of components within FL to highlight the advantages of its software design. Specifically, we utilize Appfl to evaluate the communication efficiency of different protocols, data transfer methods, and model compression algorithms. We also explore the impacts of privacy preservation algorithms on the performance of FL-trained models, as well as the training efficiency and resource utilization of various FL strategies with different synchronicity settings. Note that this section does not directly compare Appfl with other frameworks on the capabilities to resolve different FL challenges, as such framework’s capability stems from its design, consisting of a suite of components and comprehensive algorithmic and technical solutions to address diverse FL challenges. Table I already summarizes and compares those features among existing frameworks, and these solutions are generally framework-agnostic, provided a framework is capable of supporting them.
@@ IV-A Communication Efficiency
We evaluate the communication efficiency for different communication and data transfer protocols across different numbers of clients and various model sizes. Table II details the sizes of all models used in our experiments.
@@
Model # Params Size
@@
CNN 1.20M 4.58 MB
ResNet18 11.17M 42.66 MB
ResNet50 23.52M 89.93 MB
ResNet101 42.51M 162.58 MB
Vision Transformer 88.22M 336.55 MB
In the experiments, each FL client runs on a single core of a CPU node that contains two 64-core AMD EPYC 7763 “Milan” CPUs with PCIe Gen4 interfaces and 256 GB of RAM. The node is connected to the NPCF core router and exit infrastructure via two 100 gigabits per second (Gbps) connections. The FL server is hosted on an AWS EC2 x2iedn.2xlarge instance, equipped with 8 virtual CPUs, 256 GB of RAM, and up to 25 Gbps connections. We exponentially increase the number of clients from 2 to 128 across all models, except for the ViT model, which scales only from 2 to 64 because of memory constraints on the client and server hardware. We deploy all FL clients on the same hardware node to avoid performance variability caused by slower or heterogeneous machines, ensuring consistent and reproducible experimental results and accurately isolating the impact of communication protocols. We evaluate gRPC and Globus Compute communication protocols as well as two data transfer methods, AWS S3 buckets and ProxyStore endpoints. Because of the 10 MB data transfer limit with Globus Compute, it is integrated with the other two data transfer protocols rather than being tested in isolation, resulting in five distinct communication pattern combinations.
Figure 5 shows the epoch-wise average two-way communication time in seconds for various models using different communication and data transfer protocols. From the plots, we note the following key points: (1) Separating the transmission of data (i.e. model parameters) from task control signals helps communication protocols exceed their maximum data size limitations. (2) Globus Compute consistently incurs longer overheads than gRPC in transmitting control signals, which is a significant factor when the FL model size is small. (3) While data transfer via ProxyStore endpoints generally results in longer communication times, it offers a free and straightforward solution for protocols such as Globus Compute that have message size restrictions. (4) Data transfer through S3 features relatively low latency and also provides a secure, reliable means to store model checkpoints during training, although it incurs some additional costs on AWS.
@@ IV-B Compression Efficiency
We assess the efficiency of various data compression algorithms integrated within Appfl. Specifically, we utilize the lossless compressor blosc for tensors with less than 1,024 parameters and lossy compressors SZ2 [50], SZ3 [51], and ZFP [52], each with a relative error bound of 0.01, for larger tensors. The experiments adhere to the same hardware configurations described in Subsection IV-A. We conduct experiments on ResNet-50 and ResNet-101 models, scaling client numbers from 2 to 128. Figure 6(a) illustrates the reduction in model sizes by 3 to 5 times using different lossy compressors. Notably, previous studies have shown that such levels of lossy compression can preserve model accuracy within a 0.5% margin of uncompressed results [31]. Figure 6(b) presents the two-way communication times via gRPC for the two models using various compressors. Solid lines represent times with compression and decompression overheads, whereas dotted lines depict times without. The comparison reveals significant overhead, particularly with SZ2 and SZ3. Despite this, the use of compressors notably reduces communication costs and overall two-way communication times, even under high-bandwidth conditions for both clients and the server.
@@ IV-C Privacy Preservation
Model Cox model Logistic Reg. 3D U-Net EfficientNet
Metric C-index Accuracy DICE Balanced Acc.
@@ IV-D Addressing Heterogeneous Clients
Figure 8: (a) Average validation accuracy and the corresponding standard deviation on the partitioned CIFAR-10 dataset for different FL algorithms during the training process. (b) Client resource utilization for algorithms using different schedulers and the average training time per batch for different clients. (c) Visualization of computing resource utilization for three clients under different schedulers, where the colored bar represents the computing period and the blank places mean idle times.
@@
In this subsection we evaluate the performance and efficiency of different FL algorithms under various synchronicity settings. Specifically, we benchmark five FL algorithms: (1) FedAvg, a widely used synchronous algorithm that updates the global model by averaging all client local models; (2) FedAvgM, another synchronous algorithm, which incorporates momentum on top of FedAvg; (3) FedAsync, which asynchronously updates the global model upon receipt of any local model; (4) FedBuff, which is similar to FedAsync but buffers multiple local models before updating the global model; and (5) FedCompass, which introduces a COMputing Power AwarenesS Scheduler (Compass) that dynamically adjusts the number of client local training steps based on real-time estimates of client computing power to synchronize the training completion for groups of clients. As for the datasets, we partition the CIFAR-10 dataset, one of the most commonly used datasets in evaluating FL algorithms, into ten client splits in a non-IID manner, with each client holding data from five to seven classes out of ten classes. All clients use Nvidia A100 GPUs for training, and we simulate a group of heterogeneous clients by assigning different average batch processing times from an exponential distribution.
@@
Figure 8(a) presents the average validation accuracy and the corresponding standard deviation across five independent runs for each FL algorithm during training. Key observations from the figure include the following. (1) Asynchronous FL algorithms like FedAsync and FedBuff, which use the vanilla asynchronous scheduler, converge to significantly lower global model accuracy compared with synchronous methods, primarily due to the drifting toward faster clients, as the global model gets more updates from faster clients and slower clients’ models become stale. (2) Synchronous algorithms exhibit slower convergence as the server has to wait for the slow clients for aggregation. (3) FedCompass effectively addresses substantial client drift issues and attains high global model accuracy by ensuring nearly simultaneous model arrivals for grouped aggregation. It also achieves quicker convergence than synchronous methods without extensive waiting. To our best knowledge, no existing FL frameworks seamlessly support advanced FL scheduling algorithms such as FedCompass without architectural modifications, underscoring the extensibility of the Appfl framework.
@@
Figure 8(b) shows the average training time per batch for the ten clients involved in the FL training, as well as the resource utilization, calculated as the ratio of client compute time to total training time, for algorithms using the synchronous, vanilla asynchronous, and Compass asynchronous scheduler. The synchronous scheduler shows the lowest client resource utilization, correlating with training time per batch: the quicker the client, the lower the utilization. In contrast, the vanilla asynchronous scheduler, which immediately sends any received local model for aggregation and returns the updated global model, allows client resource utilization to approach 100%. Despite full utilization, however, this method results in poorly performing models due to client drift. The Compass scheduler, by estimating client speeds and adjusting training steps accordingly, maintains approximately 90% resource utilization and reduces client drift through timely grouped aggregations. Figure 8(c) visualizes the resource utilization for three clients under different scheduling scenarios, highlighting the significant resource underutilization of the synchronous scheduler compared with the asynchronous alternatives when client computing resources vary widely.
@@ V-A Vertical Federated Learning
Vertical federated learning (VFL) is a specialized paradigm of FL where different clients hold distinct features from the same dataset [44]. Unlike traditional FL (i.e., horizontal FL) dealing with the same feature space across diverse data samples, VFL enables collaboration among clients that have partially overlapping or non-overlapping features but share the same sample IDs, as illustrated in Figure 9(a). Figure 9(b) depicts a typical VFL process. In VFL, rather than training the same model architecture, each client possesses its embedding model to process its local data sample features and then sends their embeddings to the server. The server, holding the labels of the client data samples, concatenates the received embeddings to train a central model. It then sends the gradients of the feature embeddings back to the corresponding clients, enabling them to update their local embedding models accordingly.
Appfl seamlessly supports VFL by providing the VFL trainer and aggregator in the corresponding modules. In this case study, we use the diabetes datasets from the scikit-learn library, which contains ten features of 442 data samples. The labels, ranging from 25 to 346, are the responses of interest that quantitatively measure the disease progression. We split the dataset into 80% for training and 20% for validation and use three VFL clients, where clients 1 and 2 possess three patient features and client 3 possesses four. Each of the three clients as well as the server employs a two-layer perceptron with ReLU nonlinear activation as their embedding models. Figure 10(a) presents the input, hidden, and output dimensions of these models. During the training, the server model is updated based on the mean squared error (MSE) loss between the labels and predictions, using the Adam optimizer with a learning rate of 0.01. Figure 10(b) shows the training and validation MSE throughout the training.
@@ V-B Hierarchical Federated Learning
Hierarchical federated learning (HierFL) is also a special type of FL that introduces an additional role, the intermediate server (edge server). This server first aggregates local model parameters from connected clients or child intermediate servers and then forwards the aggregated model to the parent server for further aggregation [45]. HierFL is particularly beneficial when FL clients are geographically clustered, since placing an intermediate server for these clusters can significantly improve overall communication efficiency. To support HierFL in Appfl, in addition to the general server agent for the root server and the client agent for the clients, we define an intermediate server agent, inherited from the server agent, which handles FL-related requests from connected clients or child intermediate servers by interacting with its parent server.
This case study conducts four-tier HierFL experiments involving nine clients, five intermediate servers, and one root server. The MNIST dataset is partitioned into nine heterogeneous splits, with each client containing training data for only 3-5 classes. Figure 11(a) illustrates the topology of the experiment and the client data distribution. Training is conducted over 20 global epochs, with each client performing 100 local steps per epoch using a batch size of 64 and the Adam optimizer with a learning rate of 0.001. The experiments are repeated five times. Figure 11(b) presents the average validation accuracy and standard deviation for both the server model and each client’s local model on the MNIST validation set. We note that the client models are evaluated after local training. Since each client has data for only three to five classes, their local models perform significantly worse than the global model, highlighting the advantages of federated learning in leveraging data from distributed clients to train a more robust ML model.
Figure 11: (a) Topology of the multi-layer HierFL experiments. (b) HierFL validation accuracy for the server and client models, where the accuracy of client models is evaluated after each local training round.